# Tackling Graph Oversquashing by Global and Local Non-Dissipativity

### v1

@@ 3.2 Graph-wise Analysis of SWAN
While the node perspective analysis is important because it shows the long-term memory capacity of individual nodes, as illustrated in Figure 2(b), it overlooks the pairwise interactions between nodes, which is equivalently described by considering the properties of Section 3 with respect to the graph, and is illustrated in Figure 2(a).
As we show below, our SWAN is a globally non-dissipative architecture, and therefore it allows the constant rate of information flow and interactions between all nodes, independently of time $t$, equivalent to the network’s depth. Therefore, we deem that SWAN’s non-dissipativity behavior is beneficial to address oversquashing in MPNNs.
@@
$\frac{\partial\mathbf{X}(t)}{\partial t}=\sigma\Bigl{(}$ $\mathbf{X}(t){(\mathbf{W}-\mathbf{W}^{\top})}$
$+{(\hat{\mathbf{A}}+\hat{\mathbf{A}}^{\top})\mathbf{X}(t)(\mathbf{V}-{\bf V}^{\top})}$
@@
We therefore turn to analyze the ODE in Section 3.2 from a graph perspective, using the sensitivity analysis introduced in Section 2.3. Applying the vectorization operator, the Jacobian $\mathbf{J}(t)=\mathbf{M}_{1}\mathbf{M}_{2}$ writes as the multiplication of:
$\mathbf{M}_{1}$ $=\mathrm{diag}\Bigl{[}\mathrm{vec}\Bigl{(}\sigma^{\prime}\Bigl{(}\mathbf{I}\mathbf{X}(\mathbf{W}-\mathbf{W}^{\top})+$
@@
$\mathbf{M}_{2}$ $=(\mathbf{W}-\mathbf{W}^{\top})^{\top}\otimes{\bf I}+$
@@
where ${\bf I}\in\mathbb{R}^{n\times n}$ is the identity matrix, $\rm{vec}$ is the vectorization operator, and $\otimes$ is the Kronecker product (see Appendix H for more details).
Similar to our node-wise analysis in Section 3.1, $\mathbf{M}_{1}$ in Section 3.2 is a diagonal matrix, and therefore to obtain stability and non-dissipativity, we need to demand that ${\bf M}_{2}$ from Section 3.2 has eigenvalues with real part equal to zero. Indeed, we see that ${\bf M}_{2}$ satisfies this condition because it is composed of a summation of three anti-symmetric matrices, whose eigenvalues are all purely imaginary, i.e., have a real part of zero. Thus, we conclude that SWAN (Section 3.2) is both stable and globally and locally non-dissipative.
We now show that the properties of global stability and non-dissipativity allow us, theoretically, to design GNNs that can mitigate oversquashing.
@@
$\left\|\frac{\partial\rm{vec}(\mathbf{X}(t))}{\partial\rm{vec}\mathbf{X}(0))}\right\|=c,$ (13)
@@
We follow the assumption in Chang et al. (2019); Gravina et al. (2023) that the Jacobian, $\mathbf{J}(t)$, does not change significantly over time, then we can apply results from autonomous differential equations and solve Equation 14:
@@
where $\mathbf{\Lambda}$ is the diagonal matrix whose non-zero entries contain the eigenvalues of $\mathbf{J}$, and $\mathbf{T}$ has the eigenvectors of $\mathbf{J}$ as columns, as in Section 2.3. As we have previously shown, it holds that $Re(\lambda_{i}(\mathbf{J}(t)))=0$ for $i=1,...,d$, since the Jacobian is the result of the multiplication between a diagonal matrix and an antisymmetric matrix. Therefore, the magnitude of $\partial\mathbf{X}(t)/\partial\mathbf{X}(0)$ is constant over time, and the input graph information is effectively propagated through the successive layers into the final nodes’ features. ∎
The outcome of Theorem 3.1 is that regardless of the integration time $t$ (equivalent to $\ell=t/\epsilon$ layers of SWAN, where $\epsilon$ is the step size), the information between nodes will continue to propagate at a constant rate, unlike diffusion-based GNNs that exhibit an exponential decay in the propagation rate with respect to time, as shown below.
@@
A diffusion-based network with Jacobian eigenvalues with magnitude ${\bf K}_{ii}=|\mathbf{\Lambda}_{ii}|\ ,\ i\in\{0,\ldots,n-1\}$ has an exponentially decaying information propagation rate, as follows:
@@
See proof in Appendix B. This result means that with diffusion methods, after sufficiently large integration time $t$, i.e., layers, they will not be able to effectively share new information between nodes, as in early layers of the network. On the contrary, our SWAN maintains the same capacity, independent of time, meaning it retains its ability to share information across nodes with the same effectiveness in each layer of the network, regardless of its depth. We illustrate this discussion in Figure 1 and Figure 2.
@@ 3.3 The Benefit of Spatial Antisymmetry
While oversquashig was not mathematically defined in (Alon & Yahav, 2021), it was recently proposed in (Topping et al., 2022; Di Giovanni et al., 2023) to quantify the level, or lack of oversquashing, by measuring the sensitivity of node embedding after $\ell$ layers with respect to the input of another node $\left\|\partial\mathbf{x}_{v}(\ell)/\partial\mathbf{x}_{u}(0)\right\|$. Furthermore, they bounded this sensitivity score on a MPNN (Gilmer et al., 2017) as:
$\left\|\frac{\partial\mathbf{x}_{v}(\ell)}{\partial\mathbf{x}_{u}(0)}\right\|\leq\underbrace{(c_{\sigma}wp)^{\ell}}_{model}\underbrace{(\mathbf{O}^{\ell})_{vu}}_{topology}$ (17)
where $c_{\sigma}$ is the Lipschitz constant of non linearity $\sigma$, $w$ is the maximal entry-value over all weight matrices, and $p$ is the embedding dimension. The term $\mathbf{O}=c_{r}\mathbf{I}+c_{a}\mathbf{A}\in\mathbb{R}^{n\times n}$ is the message passing matrix adopted by the MPNN, with $c_{r}$ and $c_{a}$ being the weighted contributions of the residual and aggregation term, respectively. Oversquashing occurs if the right-hand side of Equation 17 is too small.
The sensitivity of SWAN. In light of the discussion above, we now provide a theoretical bound of our SWAN, with its proof in Appendix C.
@@ 4 Experiments
Objectives. We empirically evaluate our SWAN and compare it with literature methods. Specifically, we seek to address the following questions: (i) Can SWAN effectively propagate information to distant nodes? (ii) Can SWAN accurately predict graph properties that are related to long-range interactions? (iii) How does SWAN perform on real-world long-range benchmarks?
Baselines. In our experiments, the performance of SWAN is compared with various GNN baselines, namely: (i) MPNNs with linear complexity with respect to the size of nodes and edges, similarly to the complexity of our SWAN, i.e., GCN (Kipf & Welling, 2017), GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018), GatedGCN (Bresson & Laurent, 2018), GIN (Xu et al., 2019), GINE (Hu et al., 2020), and GCNII (Ming Chen et al., 2020). (ii) DE-GNNs, as a direct comparison to the family of GNNs considered in this work, such as DGC (Wang et al., 2021), GRAND (Chamberlain et al., 2021b), GraphCON (Rusch et al., 2022), and ADGN (Gravina et al., 2023). (iii) Oversquashing designated techniques such as Graph Transformers and Higher-Order GNNs, i.e., Transformer (Vaswani et al., 2017; Dwivedi & Bresson, 2021), DIGL (Gasteiger et al., 2019), MixHop (Abu-El-Haija et al., 2019), SAN (Kreuzer et al., 2021), GraphGPS (Rampášek et al., 2022), and DRew (Gutteridge et al., 2023).
Experimental Details. We provide an extensive discussion about the datasets used in our experiments in Appendix D. We discuss the complexity of SWAN and provide runtimes in Appendix F. Our method is implemented using PyTorch (Paszke et al., 2017) and PyTorch-Geometric (Fey & Lenssen, 2019), and we will openly release the code to reproduce our empirical analysis upon acceptance.
SWAN Variants. In the following experiments, we leverage the general formulation of our method and explore two main variants of SWAN, each distinguished by the implementation of the aggregation terms $\hat{{\bf A}},\ \tilde{{\bf A}}$ in the functions $\Psi$ and $\Phi$, as shown in Equation 6. Specifically, we consider
  1. (1)
    SWAN which implements the aggregation terms using pre-defined operators, which are the symmetric normalized and random walk adjacency matrices, as described in Appendix A.2.
  2. (2)
    SWAN-learn which utilizes the learned aggregation terms described in Appendix A.2.
Both variants follow our parameterization from Equation 6, and therefore are in line with the theoretical analysis provided in Section 3, and in particular Theorem 3.1. We refer the reader to Appendix E.1 for a detailed description.
@@ 4.1 Graph Transfer
Setup. We consider three graph transfer tasks, whose goal is to transfer a label from a source node to a target node, placed at distance $k$ in the graph. Due to oversquashing and dissipative behavior of existing GNN methods that are based on diffusion, the performance is expected to degrade as $k$ increases. We initialize nodes with a random valued feature and we assign values “1” and “0” to source and target nodes, respectively. We consider three graph distributions, i.e., line, ring, crossed-ring, as illustrated in Figure 4. We consider four different distances $k=\{3,5,10,50\}$. By increasing $k$, we increase the complexity of the task and require long-range information. In Appendix D.1 and E.2 we provide additional details about the dataset and the task.
Results. Figure 3 reports the results of the transfer graph tasks. The ring-graph proved to be the hardest, leading to higher errors when compared with other graphs. Overall, baselines struggle to accurately transfer the information through the graph, especially when the distance is high, i.e., #$\#hops\geq 10$. Differently, non-dissipative methods, such as ADGN and SWAN, achieve low errors across all distances. Moreover, SWAN consistently outperforms ADGN, empirically supporting our theoretical findings that SWAN can better propagate information among distant nodes.
@@
Figure 3: Information transfer on (a) Line, (b) Ring, and (c) Crossed-Ring graphs. While baseline approaches struggle to transfer the information accurately, non-dissipative methods like ADGN and our SWAN achieve low errors.
@@ 4.2 Graph Property Prediction
Setup. We address tasks involving the prediction of three graph properties - Diameter, Single-Source Shortest Paths (SSSP), and node Eccentricity on synthetic graphs as introduced in (Corso et al., 2020), and follow the experimental setup outlined in (Gravina et al., 2023), and further discussed in Appendix D.2 and E.3.
Results. In Table 1 we compare SWAN and SWAN-learn with other MPNNs and DE-GNNs. Our results indicate that SWAN improves performance across all baselines. Specifically, SWAN-learn yields the lowest (best) results across all models, showing an improvement of up to 117% with respect to the runner up model. We note that similarly to the transfer task in Section 4.1, solving Graph Property Prediction tasks necessitates capturing long-term dependencies. Hence, successful prediction requires the mitigation of oversquashing. For instance, in the eccentricity task, the goal is to calculate the maximal shortest path between a node $u$ and all other nodes. Thus, it is crucial to propagate information not only from the direct neighborhood of $u$ but also from nodes that are considerably distant.
@@
SWAN-learn -0.5981±0.1145 -3.5425±0.0830 -0.0739±0.219
@@ 4.3 Long-Range Graph Benchmark
Setup. We follow the settings in Dwivedi et al. (2022) and test SWAN on the Peptides-func and Peptides-struct datasets from the Long-Range Graph Benchmark (LRGB), on which we elaborate in Appendix D.3 and E.4. As metrics, we consider the average precision (AP) for the Peptides-func, and mean-absolute-error (MAE) for Peptides-struct.
[…]
Table 2: Performance of standard MPNNs, graph Transformers, DE-GNNs, and our SWAN across two LRGB tasks. Results are averaged over 3 weight initializations. The first, second, and third best results for each task are color-coded.
@@
GCN 0.5930±0.0023 0.3496±0.0013 0.1268±0.0060
GINE 0.5498±0.0079 0.3547±0.0045 0.1265±0.0076
GCNII 0.5543±0.0078 0.3471±0.0010 0.1698±0.0080
GatedGCN 0.5864±0.0077 0.3420±0.0013 0.2873±0.0219
@@
Transformer+LapPE 0.6326±0.0126 0.2529±0.0016 0.2694±0.0098
SAN+LapPE 0.6384±0.0121 0.2683±0.0043 0.3230±0.0039
GraphGPS+LapPE 0.6535±0.0041 0.2500±0.0005 0.3748±0.0109
@@
GRAND 0.5789±0.0062 0.3418±0.0015 0.1918±0.0097
GraphCON 0.6022±0.0068 0.2778±0.0018 0.2108±0.0091
ADGN 0.5975±0.0044 0.2874±0.0021 0.2349±0.0054
@@
SWAN 0.6313±0.0046 0.2571±0.0018 0.2796±0.0048
SWAN-learn 0.6751±0.0039 0.2485±0.0009 0.3192±0.0250

### v2

@@ 3.2 Graph-wise Analysis of SWAN
While the node-perspective analysis is important because it shows the long-term memory capacity of individual nodes, as illustrated in Figure 2(b), it overlooks pairwise node interactions, which are described by the properties of Section 3 with respect to the graph, illustrated in Figure 2(a). As we show below, our SWAN is globally non-dissipative. Hence, it allows the constant rate of information flow and node interactions, independently of time $t$, i.e., the network’s depth. Therefore, we deem that SWAN’s non-dissipativity behavior is beneficial in addressing oversquashing in MPNNs.
@@
$\frac{\partial\mathbf{X}(t)}{\partial t}=\sigma\Bigl{(}\mathbf{X}(t)$ ${(\mathbf{W}-\mathbf{W}^{\top})}+{(\hat{\mathbf{A}}+\hat{\mathbf{A}}^{\top})\mathbf{X}(t)(\mathbf{V}-{\bf V}^{\top})}$
@@
Following the sensitivity analysis introduced in Section 2.3, and applying the vectorization operator, the Jacobian of Section 3.2, $\mathbf{J}(t)=\mathbf{M}_{1}\mathbf{M}_{2}$, writes as the multiplication of:
$\mathbf{M}_{1}$ $=\mathrm{diag}\Bigl{[}\mathrm{vec}\Bigl{(}\sigma^{\prime}\Bigl{(}\mathbf{I}\mathbf{X}(\mathbf{W}-\mathbf{W}^{\top})$
@@
$\mathbf{M}_{2}$ $=(\mathbf{W}-\mathbf{W}^{\top})^{\top}\otimes{\bf I}$
@@
where ${\bf I}\in\mathbb{R}^{n\times n}$ is the identity matrix, $\rm{vec}$ is the vectorization operator, and $\otimes$ is the Kronecker product (see Appendix H for more details). As in our node-wise analysis in Section 3.1, $\mathbf{M}_{1}$ in Section 3.2 is a diagonal matrix. Thus, stability and non-dissipativity demand that ${\bf M}_{2}$ from Section 3.2 has eigenvalues with real part equal to zero. We see that ${\bf M}_{2}$ satisfies this condition as it is composed of a summation of three antisymmetric matrices whose eigenvalues have a real part of zero. We conclude that SWAN (Section 3.2) is stable, and globally and locally non-dissipative. We now show that the properties of stability and global non-dissipativity allow the design of GNNs that can mitigate oversquashing.
@@
$\left\|\frac{\partial\rm{vec}(\mathbf{X}(t))}{\partial\rm{vec}(\mathbf{X}(0))}\right\|=c,$ (14)
@@
We follow the assumption in (Chang et al. 2019; Gravina, Bacciu, and Gallicchio 2023) that the Jacobian, $\mathbf{J}(t)$, does not change significantly over time, then we can apply results from autonomous differential equations and solve Equation 15:
@@
where $\mathbf{\Lambda}$ is the diagonal matrix whose non-zero entries contain the eigenvalues of $\mathbf{J}$, and $\mathbf{T}$ has the eigenvectors of $\mathbf{J}$ as columns, as in Section 2.3. As previously shown, it holds that $Re(\lambda_{i}(\mathbf{J}(t)))=0$ for $i=1,...,d$, since the Jacobian is the result of the multiplication between a diagonal matrix and an antisymmetric matrix. Thus, the magnitude of $\partial\mathbf{X}(t)/\partial\mathbf{X}(0)$ is constant over time, allowing input features to propagate through the layers into the final node features. ∎
Theorem 3.1 states that regardless of time $t$ (equivalent to $\ell=t/\epsilon$ layers of SWAN, where $\epsilon$ is the step size), the information between nodes continues to propagate at a constant rate, unlike diffusion GNNs that exhibit an exponential decay in the propagation rate with respect to time, as shown below.
@@
A diffusion GNN with Jacobian eigenvalues with magnitude ${\bf K}_{ii}=|\mathbf{\Lambda}_{ii}|\ ,\ i\in\{0,\ldots,n-1\}$ has an exponentially decaying information propagation rate, as follows:
@@
See proof in Appendix B. This result means that for diffusion methods, as $t$ grows, deeper layers are not able to share new information between nodes as effectively as earlier layers of the network. On the contrary, our SWAN maintains the same effectiveness, independently of time, meaning it retains its ability to share information across nodes with the same effectiveness in each layer of the network, regardless of its depth, as illustrated in Figure 1 and Figure 2.
@@ 3.3 The Benefit of Spatial Antisymmetry
While oversquashig was not mathematically defined in (Alon and Yahav 2021), it was recently proposed in (Topping et al. 2022; Di Giovanni et al. 2023) to quantify the level, or lack of oversquashing, by measuring the sensitivity of node embedding after $\ell$ layers with respect to the input of another node $\left\|\partial\mathbf{x}_{v}(\ell)/\partial\mathbf{x}_{u}(0)\right\|$, which can be bounded as follows:
$\left\|\frac{\partial\mathbf{x}_{v}(\ell)}{\partial\mathbf{x}_{u}(0)}\right\|\leq\underbrace{(c_{\sigma}wp)^{\ell}}_{model}\underbrace{(\mathbf{O}^{\ell})_{vu}}_{topology},$ (18)
where $c_{\sigma}$ is the Lipschitz constant of the activation $\sigma$, $w$ is the maximal entry-value over all weight matrices, and $p$ is the embedding dimension. The term $\mathbf{O}=c_{r}\mathbf{I}+c_{a}\mathbf{A}\in\mathbb{R}^{n\times n}$ is the message passing matrix adopted by the MPNN, with $c_{r}$ and $c_{a}$ are the contributions of the residual and aggregation term. Oversquashing occurs if the right-hand side of Equation 18 is too small (Di Giovanni et al. 2023).
The sensitivity of SWAN. We present the sensitivity bound of SWAN, with its proof in Appendix C.
@@ 4 Experiments
Objectives. We evaluate our SWAN and compare it with various methods. Specifically, we seek to address the following questions: (i) Can SWAN effectively propagate information to distant nodes? (ii) Can SWAN accurately predict graph properties related to long-range interactions? (iii) How does SWAN perform on real-world long-range benchmarks?
Baselines. We consider (i) MPNNs with linear complexity (similar to the complexity of our SWAN), i.e., GCN (Kipf and Welling 2017), GraphSAGE (Hamilton, Ying, and Leskovec 2017), GAT (Veličković et al. 2018), GatedGCN (Bresson and Laurent 2018), GIN (Xu et al. 2019), GINE (Hu et al. 2020b), and GCNII (Chen et al. 2020). (ii) DE-GNNs such as DGC (Wang et al. 2021), GRAND (Chamberlain et al. 2021b), GraphCON (Rusch et al. 2022), and ADGN (Gravina, Bacciu, and Gallicchio 2023). (iii) Oversquashing designated methods such as Graph Transformers and Higher-Order GNNs, i.e., Transformer (Vaswani et al. 2017; Dwivedi and Bresson 2021), DIGL (Gasteiger, Weißenberger, and Günnemann 2019), MixHop (Abu-El-Haija et al. 2019), SAN (Kreuzer et al. 2021), GraphGPS (Rampášek et al. 2022), and DRew (Gutteridge et al. 2023).
Experimental Details. We discuss the datasets in our experiments in Appendix D, and the complexity of SWAN and provide runtimes in Appendix F. Our implementation uses PyTorch, and is available at https://github.com/gravins/SWAN.
SWAN Variants. In the following experiments, we leverage the general formulation of our method, discretized by forward Euler (see Appendix A.1), and explore two main variants of SWAN, each distinguished by the implementation of the aggregation terms $\hat{{\bf A}},\ \tilde{{\bf A}}$ in the functions $\Psi$ and $\Phi$, as shown in Equations 6 and 7. Specifically, we consider (i) SWAN, which implements the aggregation terms using pre-defined operators, which are the symmetric normalized and random walk adjacency matrices, as described in Appendix A.2, and, (ii) SWAN-learn which utilizes the learned aggregation terms described in Appendix A.2. Both variants follow the form of Equations 6 and 7, in line with the theoretical analysis in Section 3, and in particular Theorem 3.1. We refer the reader to Appendix E.1 for a detailed description.
@@ 4.1 Graph Transfer
Setup. We consider a graph transfer task, where the goal is to transfer a label from a source to a target node, with a distance of $k$ hops. We note that this task can be effectively solved only by non-dissipative methods that preserve source information. We initialize nodes with a random valued feature and we assign values “1” and “0” to source and target nodes, respectively. We consider three graph distributions, i.e., line, ring, crossed-ring, as illustrated in Figure 4, with four different distances $k=\{3,5,10,50\}$. Increasing $k$ increases the task complexity. In Appendix D.1 and E.2, we provide additional details about the dataset and the task.
Results. Figure 3 reports the results on graph transfer tasks. Overall, baseline methods struggle to accurately transfer the information through the graph, especially when the distance is high, i.e., #$\#hops\geq 10$. Differently, non-dissipative methods, such as ADGN and SWAN, achieve low errors across all distances. Moreover, SWAN consistently outperforms ADGN, empirically supporting our theoretical findings that SWAN can better propagate information among distant nodes.
@@
Figure 3: Information transfer performance on (a) Line, (b) Ring, and (c) Crossed-Ring graphs. Non-dissipative methods like ADGN and SWAN allow for the accurate transfer of information.
@@ 4.2 Graph Property Prediction
  SWAN-learn -0.5981±0.1145 -3.5425±0.0830 -0.0739±0.2190
Setup. We consider the prediction of three graph properties - Diameter, Single-Source Shortest Paths (SSSP), and node Eccentricity on synthetic graphs (Corso et al. 2020), and follow the setup outlined in Gravina, Bacciu, and Gallicchio (2023), further discussed in Appendix D.2 and E.3. These three tasks rely on long-range interactions for shortest-path calculations, as in Bellman-Ford and Dijkstra’s algorithms.
Results. In Table 1, we compare SWAN and SWAN-learn with MPNNs and DE-GNNs, indicating that SWAN consistently improves performance, where SWAN-learn yields the best results, with an improvement of up to 117% of the runner-up model. In Appendix G we also provide results for SWAN with layer-dependent weights showing improved performance. We note that similarly to the transfer task in Section 4.1, solving Graph Property Prediction tasks necessitates capturing long-term dependencies. Hence, successful prediction requires the mitigation of oversquashing. For instance, in the eccentricity task, the goal is to calculate the maximal shortest path between node $u$ and all other nodes – a task requiring propagation from distant nodes.
@@ 4.3 Long-Range Graph Benchmark
Setup. We follow the settings from Dwivedi et al. (2022) on the Peptides-func, Peptides-struct, and PascalVOC-sp datasets, on which we elaborate in Appendix D.3 and E.4. As metrics, we consider the average precision (AP) for the Peptides-func, mean-absolute-error (MAE) for Peptides-struct, and the macro-weighted F1 score for PascalVOC-sp.
Results. Table 2 compares SWAN with various MPNNs, DE-GNNs, and Graph Transformers.
@@
  GCN 59.30±0.23 0.3496±0.0013 12.68±0.60
  GINE 54.98±0.79 0.3547±0.0045 12.65±0.76
  GCNII 55.43±0.78 0.3471±0.0010 16.98±0.80
  GatedGCN 58.64±0.77 0.3420±0.0013 28.73±2.19
@@
  Transformer+LapPE 63.26±1.26 0.2529±0.0016 26.94±0.98
  SAN+LapPE 63.84±1.21 0.2683±0.0043 32.30±0.39
  GraphGPS+LapPE 65.35±0.41 0.2500±0.0005 37.48±1.09
@@
  GRAND 57.89±0.62 0.3418±0.0015 19.18±0.97
  GraphCON 60.22±0.68 0.2778±0.0018 21.08±0.91
  ADGN 59.75±0.44 0.2874±0.0021 23.49±0.54
@@
  SWAN 63.13±0.46 0.2571±0.0018 27.96±0.48
  SWAN-learn 67.51±0.39 0.2485±0.0009 31.92±2.50
Table 2: Performance of standard MPNNs, graph Transformers, DE-GNNs, and our SWAN across three LRGB tasks. Results are averaged over 3 weight initializations. The first, second, and third best results for each task are color-coded.
[…]
Dataset $\downarrow$ / Model $\rightarrow$ SWANβ=0 SWAN-ne SWAN-ne-learn SWAN SWAN-learn
Diam. ($log_{10}$(MSE)$\downarrow$) -0.3882±0.0610 -0.5497±0.0766 -0.5631±0.0694 -0.5249±0.0155 -0.5981±0.1145
SSSP ($log_{10}$(MSE)$\downarrow$) -3.2061±0.0416 -3.1913±0.0762 -3.5296±0.0831 -3.2370±0.0834 -3.5425±0.0830
Ecc. ($log_{10}$(MSE)$\downarrow$) 0.5573±0.0247 0.3792±0.1514 0.1317±0.1253 0.4094±0.0764 -0.0739±0.219
Peptides-func (AP $\uparrow$) 61.95±00.67 61.19±0.37 62.49±0.51 63.13±0.46 67.51±0.39
Peptides-struct (MAE $\downarrow$) 0.2703±0.0023 0.2672±0.0012 0.2606±0.0007 0.2571±0.0018 0.2485±0.0009
Table 3: Performance of different versions of SWAN on Graph Property Prediction and LRGB tasks. Results are averaged over 4 random weight initializations on the Graph Property Prediction, while over 3 on the LRGB. The first, second, and third best results for each task are color-coded. Global and Local Non-Dissipative variants achieve highest performance.